Solve the game with the payoff Matrix 2 6 -4 -7 4 7 -3 -5 3 3 9 8
Answer #1

Maximim for A=Max(Minimum in each Row)

=Max(-7,-5,3)

=3

Minimax of B=Min(Maximum in each column)

=Min(4,7,9,8)

=4

Maximum for A is called the lower value of the game and denoted by V and minimax of B is called the upper value of the game and denoted by \bar{V}

So here V=3 & \bar{V} =4

If V is the value of the game then always satisfies the inequality

Maximim for A\leq V \leq minimax for B

  3\leq V \leq 4

If \bar{V} =V=V then saddle point exist

But here \bar{V} \neqV

so, here saddle point is not exist.

Thank you ,

If any queries please leave your valuable comment on comment box

Please rate my answer as well